Hook Industries's capital structure consists solely of debt and common equity. It can issue debt at rd = 9%, and its common stock currently pays a $3.50 dividend per share (D0 = $3.50). The stock's price is currently $26.75, its dividend is expected to grow at a constant rate of 4% per year, its tax rate is 25%, and its WACC is 15.55%. What percentage of the company's capital structure consists of debt?
Answer #1

Let’s first calculate the cost of equity by using dividend discount model with constant dividend growth rate

Current Stock Price P0 = D0 * (1+g) / (r – g)

Where:

P0 = the current stock price = $26.75

Last dividend paid D0 = $3.50 per share

g = growth rate of dividends = 4% per year

re = required rate of return or cost of equity =?

Therefore

$26.75 = $3.50 * (1+4%) / (re - 4%)

Or $26.75 *(re - 4%) = $3.64

Or $26.75 *re - $26.75 *4% = $3.64

Or $26.75 *re - $1.07 = $3.64

Or $26.75 * re = $3.64 + $1.07 = $4.71

Or re =$4.71 /$26.75 = 0.1761 or 17.61%

Therefore the company’s cost of common equity is 17.61%

Now to calculate the weighted cost of capital (WACC), assume that weight of debt is X and weight of equity is (1-X)

WACC = wd *rd (1 - t) + we*re                 

Where,

WACC = 15.55%

we is the weight of equity = 1-X

wd is value of debt = X

rd is the before-tax cost of debt = 9%

t is the company’s tax rate = 25%

re is the cost of equity = 17.61% (calculated above)

Therefore,

15.55% = X * 9% * (1-25%) + (1-X) * 17.61%

Or 15.55% = 6.75% X + 17.61% - 17.61% X

Or 17.61% X -6.75% X = 17.61% - 15.55%

Or X = 2.06%/ 10.86% = 0.1895 or 18.95%

The percentage of the company's capital structure consists of debt is 18.95%